P09701: Corning Tropel LightGage Metrology System
/public/

Week 9:The Build Continues

Week 9: The Build Continues (by Friday 5/8/09)

All documents generated by the P09701 team during Week 9 of MSD2 are to be posted here.

MSD2 Week 9 Progress

Monday's News

1. Wire management implemented on air table.

2. Repeatability tests begun.

Tuesday's News

1. We have learned that actuator #3 is still in the possession of the USPS, and the status of the package is unknown. The USPS has been called and an inquiry has been opened into the case. However, only the package sender is privy to USPS information. Mirai has said that they will keep us posted if they learn anything.

2. We should still be able to get the system working with only 2 motors. It is pretty well balanced, and each motor should be able to easily handle the weight.

3. Lift pads have been finished!

4. Preliminary repeatability test ran overnight to see what kind of measurement quality we can get out of the top head. Results below for the last 25 measurements.

public/Resultsexcel.png

As can be seen, the standard deviation of the flatness is over .5 microns, which is not very good. Will ask Tropel how we might be able to make measurements better.

public/TMSresults.png

5. Actuators have been integrated into TMS! They can now be controlled through ManMotion. We should be able to lift head tomorrow.

6. Chris will come in on Wednesday with a second head to work on getting measurements with 2 heads.

Wednesday's News

1. Chris came with a second head and hooked it up. Unfortunately, the computer that was given to us is way too slow to adequately handle both streaming videos. It works, though. Albeit very slowly.

Thursday's News

1. All three motors are moving in tandem.

2. Scripts have been written to make them move, and have been integrated into TMS. Now when you push the measure button, the actuators lift the head up, prompt user to place part on harp, puts it back down, and then lifts it back up. So basically, it is working!

3. Friday will bring more refinement.

Friday's News

1. Prompts were added to the scripts to notify user of what is happening. This seems to work well.

2. Quick-reference guide worked on for customer.

Saturday's News

1. Worked on figuring out why contrast was so low on top head. Adjusted coherence disk position and fixed this problem. Now fringes are very high contrast, and the whole window shows up in every measurement. Adjusted focus as well--parts look good. There seems to be some alignment error on the bottom head (as Chris predicted when he gave it to us), so there is a shadow on the part. Aligning the optics will solve this problem.

2. Played with recipe to check all the regions are extracting as expected. Turns out there is a software glitch where TMS will not allow the user to input a height for a region--it just defaults to 169mm. Odd. Also, the + and - ambiguity buttons in the region finder both go in the + direction. Another glitch?

3. System is near 100% working. All that remains is:

a. Complete and install 3rd lift pad.

b. Get part detection sensor to work with the motors.

c. Build a recipe to test the system's capabilities (may be limited because of Tropel software glitches)

d. Wire management.

e. Clean up computer for delivery to customer.

f. Week 11 Poster.

g. Tech. paper.